Wavelength (nm) 275 KE (J*1019) 3.403 2.819 300 1/Wavelenth (nm-1) 0.003636364 0.003333333 0.003076923 0.002857143 0.00266666 Based on this information, how can I calculate (a) the work function (phi), (b) the value of Planck's constant and (c) the maximum wavelength, in nm, for the photoelectric effect to manifest?
Answer #1

Salo E W + KE(max) W + KEmas he C KE max= y, ſ = x) KE max= [-W) + he y= heat (-W) y = 2031.84+13.9847)00) So. work function 3.9947 Jx lots A 3,9847%2019 @ 19 (b) hc - 20 31.8 x -28 2031.8.x to J hm Jm. -34 he 2031.8 x do -28 3x